A college student earned $9000 during summer vacation working as a waiter. the student invested part of the money @ 8% and the rest @ 7%. if the student recieved a total of %667 in interest at the end of the year, how much was invested @ 8% ?
Answer by jorel1380(3719)

x+y=9000
.08x+.07y=667
-------------
8x+7y=66700
7x+7y=63000
x=3700
y=5300
$3700 was invested at 8%, $5300 was invested at 7%..